# Optimal rates of approximation by shallow ReLUk neural networks and applications to nonparametric regression

### 21 Jun 2023 v2

@@ 5 Conclusion
There are a few interesting questions we would like to propose for future research. First, for shallow neural networks approximation, the optimal rate $\mathcal{O}(N^{-\alpha/d})$ in the supremum norm (in Corollary 2.4) is established only for ReLU and Heaviside activations. Can this result be extended to all ReLUk activations? Second, it is unclear whether over-parameterized neural networks can achieve the optimal rate for learning functions in $\mathcal{H}^{\alpha}$. It seems that refined generalization error analysis is needed. Finally, it would be interesting to extend the theory developed in this paper to general activation functions.
@@ Acknowledgments
The work described in this paper was partially supported by InnoHK initiative, The Government of the HKSAR, Laboratory for AI-Powered Financial Technologies, the Research Grants Council of Hong Kong [Projects No. CityU 11308121, 11306220, 11308020 and No. N_CityU102/20] and National Natural Science Foundation of China [Projects No. 11461161006].

### 09 Jan 2024 v3

@@ 5 Conclusion
There are a few interesting questions we would like to propose for future research. First, for approximation by shallow neural networks, we establish the optimal rate in the supremum norm by using the results of [55] (Lemma 2.3). The paper [55] actually showed that approximation bounds similar to Lemma 2.3 also hold in Sobolev norms. We think it is a promising direction to extend our approximation results in the supremum norm (Theorem 2.1 and Corollary 2.4) to the Sobolev norms. Second, it is unclear whether over-parameterized neural networks can achieve the optimal rate for learning functions in $\mathcal{H}^{\alpha}$. It seems that refined generalization error analysis is needed. Finally, it would be interesting to extend the theory developed in this paper to general activation functions and study how the results are affected by the activation functions.
@@ Acknowledgments
The work described in this paper was partially supported by InnoHK initiative, The Government of the HKSAR, Laboratory for AI-Powered Financial Technologies, the Research Grants Council of Hong Kong [Projects No. CityU 11306220 and 11308020] and National Natural Science Foundation of China [Project No. 12371103] when the second author worked at City University of Hong Kong. We thank the referees for their helpful comments and suggestions on the paper.